OpenStudy (miss.rose):

I need someone to look over 5 questions. Can someone please help me? I will fan and medal.

MsBrains (ms-brains):

Sure!

OpenStudy (boldjon):

wooh @Ms-Brains is here

MsBrains (ms-brains):

Lol. @boldjon Yup! :P

OpenStudy (miss.rose):

Okay. This is the first question out of five: I think it is 5.

OpenStudy (mathmale):

@Miss.Rose: please show your work, so we'll know why you think the correct answer choice is 5. Thank you.

OpenStudy (miss.rose):

This is just a guess. So I don't know how to solve this. @mathmale

OpenStudy (mathmale):

I appreciate your admitting that. Let's learn how to solve "direct variation" problems. If y = k*x, then y increases when x increases, and y decreases when x decreases. This is the meaning of direct variation. Our job is to find the value of the "constant or proportionality," which is "k".

OpenStudy (mathmale):

choose an (x,y) pair from the table now, please. Any one you pick should be fine.

OpenStudy (miss.rose):

(x) 4 (y) 20

OpenStudy (mathmale):

Great. Now, substitute y=20 and x=4 into the equation y=kx. Show your work. Our goal is to solve for k.

OpenStudy (mathmale):

Re-write y=kx subbing 20 for y and 4 for x.

OpenStudy (miss.rose):

20=k4

OpenStudy (mathmale):

Great. Now, solve for k.

OpenStudy (miss.rose):

How do I solve for k?

OpenStudy (mathmale):

Isolate k. To do this, divide both sides of the equation by 4.

OpenStudy (miss.rose):

20div4=5 k4div4=k1 I am not sure, is this how I do it?

OpenStudy (mathmale):

\[\frac{ 20 }{ 4 }=k \frac{ 4 }{ 4 }\]

OpenStudy (mathmale):

Reduce that. 20/4=? 4/4=?

OpenStudy (miss.rose):

20/4=5 4/4=1

OpenStudy (mathmale):

Right. k=5. Re-write y=kx, this time subbing 5 for k.

OpenStudy (miss.rose):

20=54?

OpenStudy (mathmale):

First, Miss.Rose, type y=kx. Just copy this formula.

OpenStudy (miss.rose):

y=kx 20=5*4?

OpenStudy (mathmale):

Your y=kx is correct, as is your example. However, I would like for you to type y=kx and then type the same thing, substituting 5 for x. Please do that now.

OpenStudy (miss.rose):

y=5k

OpenStudy (mathmale):

exactly. THAT is the direct proportion relationship we wanted.

OpenStudy (mathmale):

k is the "constant of variation." What is k in your equation, above?

OpenStudy (miss.rose):

4

OpenStudy (mathmale):

How so? You got 5 before.

OpenStudy (miss.rose):

Sorry, I thought you said what is x. k=5

OpenStudy (mathmale):

Now look at the four answer choices. Which one is correct?

OpenStudy (miss.rose):

5?

OpenStudy (mathmale):

Right. Congrats.

OpenStudy (miss.rose):

Thank you, can you help me with 4 more?

OpenStudy (mathmale):

One more. My goal is to help you learn how to solve those problems yourself. Choose a hard one.

OpenStudy (miss.rose):

I am going to give you the choice of the 3 I think are hard. Look at them and then decide which one is the hardest, and then help me solve it.

OpenStudy (mathmale):

Look carefully at question #5. What do you already know about solving this type of question?

OpenStudy (miss.rose):

Nothing. I missed out on a lesson at school, and I asked my friend for her notes, it turns out her notes were useless.

OpenStudy (mathmale):

so sorry. That happens. that's why I prefer to lead you through all of the problem solving steps.

OpenStudy (miss.rose):

Thank you.

OpenStudy (mathmale):

the most important words here are "varies directly." This also describes the previous problem we worked on. We got y = 5 x, remember? k is the "constant of proportionality."

OpenStudy (miss.rose):

Yes, I remember.

OpenStudy (mathmale):

The dependent variable is C, which represents the number of Calories consumed. C varies directly with the time spent exercising.

OpenStudy (mathmale):

Again, C is your dependent variable; its value depends on the value of t. t is your independent variable; it represents how many hours are spent exercising, and varies from zero to whatever.

OpenStudy (mathmale):

k is just the constant of proportionality, which was 5 in the previous problem. Give this a stab: Write a "direct variation" equation with C as dependent variable, t as independent variable, and k as "constant of proportionality."

OpenStudy (mathmale):

By all means review our previous discussion if y ou wish.

OpenStudy (miss.rose):

C=tk

OpenStudy (mathmale):

I'd write that as C = k t, but otherwise your result is fine!

OpenStudy (miss.rose):

Okay. C=kt

OpenStudy (mathmale):

Cool

OpenStudy (mathmale):

Do both of the remaining 2 problems yourself. Msg me when you're done.

OpenStudy (miss.rose):

Wait... The answers only have C=800t... Is that what the answer is? @mathmale

OpenStudy (mathmale):

Ooops. We didn't find the value of k. Pick any one pair (t, C) from the table given.

OpenStudy (mathmale):

It says that after this guy walks 4 hours, he has burned 800 calories. Agree with that?

OpenStudy (miss.rose):

Yes. So would it be: C=800t

OpenStudy (mathmale):

Not quite. WRite C = k t. Now let C = 800 cal and t = 4 hr. Find k, the constant of proportionality.

OpenStudy (miss.rose):

C=kt 800=k4 So we divide 800 by 4 which is 200.

OpenStudy (mathmale):

so k=200. re-write C = k t, subbing 200 for k.

OpenStudy (miss.rose):

C=200t?

OpenStudy (mathmale):

right. very good! done. sorry I asked you to get info from a table; there's no table for this problem!

OpenStudy (mathmale):

Which of the four answers is the correct one?

OpenStudy (miss.rose):

C=200t

OpenStudy (mathmale):

Done.

OpenStudy (miss.rose):

Okay. I am not really sure how to solve number 3, can you help?

OpenStudy (mathmale):

Have to excuse myself now, but I'll be lurking in the background if you need me. You're doing well. Please at least try #3 on your own first. When you've done that, tell me what you know about solving it and what you're unsure of.

OpenStudy (miss.rose):

Okay.

OpenStudy (miss.rose):

So... I tried solving it myself and I got: y=2/5x I got this because I divided 10 by 4 and got 2.5 and turned that into a fraction. Did do that right, and get the correct answer.

OpenStudy (mathmale):

First, the problem states, "y varies directly with x." You should take that as an order to write the appropriate direct variation equation. We've done that several times before. Can you remember this direct variation equation?

OpenStudy (mathmale):

Recall y = k x. This is the general equation of ":direct variation." OK. Problem states: "y=10 when x=4." Can you re-write y = k x using that info? Can y ou then find the constant of variation, k?

OpenStudy (miss.rose):

y=kx

OpenStudy (mathmale):

right. you are told that y=10 when x=4. substitute these values into y = k x.

OpenStudy (miss.rose):

10=k4

OpenStudy (mathmale):

Now divide both sides of the equation by 4 to find the value of the proportionality constant, k. Please show all your work.

OpenStudy (miss.rose):

10/4=2.5 4/4=1

OpenStudy (miss.rose):

So... then I turn 2.5 into a fraction, right? \[\frac{ 2 }{ 5 }\]

OpenStudy (mathmale):

Divide 10 = k (4) by 4.... we get (10/4) = k. Note that 10/4 is greater than 1. 10/4 = 5/2. either is OK, 5/2 is better. 2/5 is the reciprocal of 5/2, and is not ok. What is your new k value?

OpenStudy (miss.rose):

5/2

OpenStudy (miss.rose):

\[y=\frac{5 }{2 }x\]

OpenStudy (mathmale):

right. k = 5/2. Now, as your last step, write the equation of direct variation in the form y = (k) x. Yes. You have it.

OpenStudy (mathmale):

congrats!

OpenStudy (miss.rose):

Thank you. How do you solve number 2? I have no clue what the first step is. Again, THANK YOU!

OpenStudy (mathmale):

Look at y = 40x. Borrowing from previous experience, what kind of variation is modeled here?

OpenStudy (miss.rose):

constant

OpenStudy (mathmale):

Actually, that 's "direct proportion" or "direct variation." Important terms. You are given y = 40x, which has the same form as y=kx, which we used earlier. In this case, k=40 is the proportionality constant.

OpenStudy (mathmale):

Now re-read Question 2. Which answer do you think is correct?

OpenStudy (miss.rose):

40?

OpenStudy (mathmale):

Right. You 'borrowed' that from y = 40 x. That's absolutely correct!

OpenStudy (miss.rose):

Thank you sooooo much for ALL of your help!!!!!!

OpenStudy (mathmale):

'Twas great working with you. I appreciate your enthusiasm. Hope to work with you again. Bye for now!